# Transmission lines;Defect detection;UAV imagery;Small Object Detection;Transformer-based detectors

### v6

@@ Abstract
Transmission lines;Defect detection;UAV imagery;Small Object Detection;Transformer-based detectors
@@ 3.1 Overview
In this study, we employ RT-DETR-r18Zhao et al. (2024) as the baseline detection framework owing to its superior accuracy and robustness in complex scenarios. Traditional DETR models often suffer from slow convergence and substantial computational overhead, which hinder their deployment in UAV-based inspection tasks.
Addressing these inherent limitations, RT-DETR incorporates an enhanced encoder–decoder architecture together with refined query interaction mechanisms, while for the backbone network, RT-DETR-r18 adopts a CNN-based design, namely the popular ResNet-18He et al. (2016). From the perspective of real-time performance, employing a CNN architecture for feature extraction further enhances the real-time capability of DETR-like models, thereby improving their practical applicability. As a result, RT-DETR not only accelerates convergence and improves detection precision, but also preserves the global context modeling capability of Transformer-based frameworks(Vaswani et al., 2017).In addition, its lightweight optimization strategies—specifically, the design choice to apply the encoder only to high-level semantic features (S5) with relatively small spatial dimensions—effectively reduce resource consumption. This balances performance and speed by significantly decreasing computational overhead and accelerating inference without compromising accuracy, thereby making RT-DETR particularly suitable for post-flight UAV imagery analysis, where large volumes of data must be processed efficiently under computational and energy constraints(Tian et al., 2025).
@@ 3.4 Cross-Stage Dual-Domain Multi-Scale Attention Module
In transmission line defect detection, most objects are small-scale objects such as cracks, rust spots, or insulator damages. These defects often exhibit weak visual saliency and subtle structural variations, making them difficult to distinguish from complex backgrounds such as vegetation, sky, or towers. Conventional convolutional networks are limited by their local receptive fields, while large kernels may cause oversmoothing and incur high computational costs. Similarly, traditional attention mechanisms focusing solely on spatial or channel domains fail to provide sufficient enhancement for ambiguous and small objects. Therefore, a more effective mechanism is required to preserve fine-grained details, capture global semantics, and adaptively highlight defect-related features.
To address these challenges, we propose the Cross-Stage Dual-Domain Multi-Scale Attention Module (CSDMAM). Given input features $X\in\mathbb{R}^{B\times C\times H\times W}$, we first employ a $1\times 1$ convolution to adjust the channel dimension. The transformed features are then fed into parallel branches: a large-receptive-field branch, a local branch, and a dual-domain Frequency–Spatial–Channel Attention (FSCA) branch. The outputs are aggregated through residual summation and projected by a $1\times 1$ convolution to form the final response.To provide a clearer understanding of the proposed design, the overall architecture of CSDMAM is illustrated in Figure 5.
Figure 5: The architecture of Cross-Stage Dual-Domain Multi-Scale Attention Module (CSDMAM).
To effectively capture different types of information, we incorporate four complementary branches into the module.
@@
To strengthen semantic discrimination, we design a Frequency–Spatial–Channel Attention (FSCA) branch that jointly exploits frequency, spatial, and channel dependencies. Given input features $X\in\mathbb{R}^{B\times C\times H\times W}$, we first apply the two-dimensional Fourier transform to obtain the frequency representation
@@
where $\mathcal{F}$ denotes the discrete Fourier transform (DFT). Channel-wise weights are then computed via global average pooling followed by a $1\times 1$ convolution, producing $A_{c}\in\mathbb{R}^{B\times C\times 1}$. The enhanced frequency map is obtained as
@@
which is reconstructed into the spatial domain using the inverse DFT:
@@
An additional spatial–channel reweighting map $A_{sc}$ is generated by global pooling and a $1\times 1$ convolution, yielding the FSCA-enhanced representation
@@
The outputs of all branches are fused with the residual input to form the aggregated response:
@@
where $DW$ denotes depth-wise convolution with the specified kernel size. Finally, to improve efficiency and gradient propagation, we extend the design into a cross-stage structure. Following the Cross Stage Partial (CSP) strategy, the input is divided into two parts: one processed by the multi-branch attention operations and the other preserved as an identity mapping. The two parts are concatenated and fused through a $1\times 1$ convolution, giving the final output
$Y_{cs}=\text{Conv}([\text{DMAM}(X_{\text{branch}}),X_{\text{identity}}]).$
By combining large receptive fields, local detail preservation, and FSCA-based dual-domain enhancement within a cross-stage framework, CSDMAM achieves a favorable balance between global semantic understanding and precise defect localization, making it particularly effective for detecting small, visually ambiguous objects in UAV inspection imagery.
@@ 3.5 Focaler-Wise-SIoU Loss Function
In the task of transmission line defect detection, where the majority of objects are small and often exhibit subtle spatial differences, accurate bounding box regression plays a crucial role. IoU-based loss functions have become the de facto standard for this purpose in modern object detectors, yet they still suffer from several inherent limitations. The vanilla IoU loss only considers the overlap ratio and fails to capture spatial misalignment when boxes do not intersect. Its variants, such as GIoU(Rezatofighi et al., 2019), DIoU(Zheng et al., 2020), and CIoU(Zheng et al., 2022), partially alleviate this issue by introducing penalties related to enclosure area, center distance, or aspect ratio. However, these approaches remain limited in their ability to jointly model geometric alignment and shape consistency. More advanced designs such as SIoU(Gevorgyan, 2022) attempt to incorporate angular, distance, and shape discrepancies, providing richer geometric information. Nevertheless, they still treat all samples equally, regardless of their difficulty, which results in an imbalanced gradient contribution. Easy samples dominate the optimization, while hard samples, particularly those corresponding to small defects with low signal-to-noise ratios, receive insufficient emphasis. Furthermore, in high-IoU regimes where precise localization is critical for tiny defect regions, conventional losses often lack discriminative power, leading to slower convergence and degraded accuracy.
To address these limitations, we propose the Focaler-Wise-SIoU Loss, which integrates normalized IoU scaling, comprehensive geometric penalties, and a non-monotonic focal modulation scheme. This loss is specifically designed to enhance sensitivity in high-IoU regimes and to dynamically reweight samples according to their difficulty, making it particularly suitable for datasets dominated by small and challenging objects such as transmission line defects.
Formally, given a predicted box $B_{p}=(x_{p},y_{p},w_{p},h_{p})$ and a object box $B_{t}=(x_{t},y_{t},w_{t},h_{t})$, both parameterized by center coordinates and width-height pairs, we first convert them into corner representation $(x_{1},y_{1},x_{2},y_{2})$. The intersection and union regions are then computed, yielding the standard Intersection over Union score:
$IoU=\frac{S_{int}}{S_{union}}.$ (5)
To improve discriminability in high-IoU ranges, we apply Focal-IoU normalization, which linearly maps IoU into the interval $[0,1]$ as:
$IoU=\text{clip}\left(\frac{|IoU-d|}{u-d},0,1\right),$ (6)
where $d=0$ and $u=0.95$. The base penalty is then defined as $I_{IoU}=1-IoU$.
On top of this formulation, we incorporate the Scylla-IoU (SIoU) penalty, which considers angle, distance, and shape discrepancies. The angular component is derived from the displacement between box centers $\Delta=(d_{x},d_{y})$ as:
@@
$\text{Angle}=\sin(2\phi)-2.$ (8)
This term modulates the distance penalty:
$\text{Dist}=2-\exp\left(\text{Angle}\cdot\frac{d_{x}}{w_{box}}\right)-\exp\left(\text{Angle}\cdot\frac{d_{y}}{h_{box}}\right),$ (9)
where $(w_{box},h_{box})$ denote the dimensions of the smallest enclosing box. To model shape consistency, we define the shape cost as:
$\text{Shape}=\left(1-\exp\left(-\frac{|w_{p}-w_{t}|}{\max(w_{p},w_{t})}\right)\right)+\left(1-\exp\left(-\frac{|h_{p}-h_{t}|}{\max(h_{p},h_{t})}\right)\right)^{\theta},$ (10)
with $\theta=4$. The SIoU core-loss is then expressed as:
$L_{SIoU}=L_{IoU}+\frac{\text{Dist}+\text{Shape}}{2}.$ (11)
To ensure adaptive treatment of samples, particularly the abundance of small and hard defect objects, we introduce a Wise modulation mechanism. A relative difficulty factor is defined as:
$\beta=\frac{L_{IoU}}{\mathbb{E}[L_{IoU}]}.$ (12)
where $\mathbb{E}[L_{IoU}]$ is an exponential moving average updated with a momentum of $10^{-2}$. Unlike conventional monotonic modulations, which continuously emphasize harder cases, we employ a non-monotonic focal modulation (V3) strategy to avoid over-penalization of extreme outliers:
$\gamma(\beta)=\frac{\beta}{\delta\cdot\alpha^{(\beta-\delta)}}.$ (13)
This modulation adaptively emphasizes moderately difficult samples while suppressing trivial and excessively noisy cases, which is particularly beneficial for stabilizing training in small-object detection.
Finally, the complete Focaler-Wise-SIoU Loss is formulated as:
$L=\left(1-\hat{IoU}+\frac{\text{Dist}+\text{Shape}}{2}\right)\cdot\frac{1-IoU}{E[1-IoU]}\cdot\frac{1-IoU}{E[1-IoU]}^{\delta}.$ (14)
This unified design preserves the interpretability of IoU-based losses while enhancing discriminative power in high-IoU regimes, adaptively balancing sample contributions, and stabilizing optimization. As a result, the Focaler-Wise-SIoU Loss is particularly effective in transmission line defect detection, where most objects are small, difficult to localize, and require precise bounding box regression.
@@ 4.1 CSG-ADCD
The CSG-ADCD (China Southern Grid Aerial Defective Component Dataset) is a large-scale, high-resolution UAV-based image dataset specifically developed for the detection of defects in power transmission line components, including insulators, tie wires, and poles. Constructed under authentic inspection scenarios by China Southern Power Grid, the dataset comprises 10,000 carefully selected aerial images. Each image was manually annotated by a single domain expert to ensure both consistency and high-quality labeling.
In total, the dataset contains 73,448 annotated instances covering nine component states or defect types: Normal Insulator (zcjyz), Polluted Insulator (jyzwh), Damaged Pole (dgss), Missing Tie Wire (zxqs), Loose Tie Wire (zxst), Insulator Flashover (jyzsl), Bird Nest (nw), Broken Insulator (jyzps), and Shattered Post Insulator (zyzsl). Among these, Normal Insulator is the dominant class with 43,808 instances, whereas Shattered Post Insulator is the rarest with only 37 instances.
@@
Table 1: The correspondence between labels and category names.
NO Label Name
@@
As shown in Fig.6, the vast majority of objects are smaller than 32×32 pixels, placing them within the small-object category. The scatter plot depicts the distribution of annotated instances by pixel width and height after preprocessing, with different categories distinguished by distinct colors in the legend. The strong concentration of points in the lower-left corner highlights the predominance of small and compact objects, consistent with the characteristics of transmission-line inspection imagery. Dashed reference lines at 32×32 and 96×96 pixels mark the boundaries between small, medium, and large regimes, clearly revealing the relative scarcity of medium- and large-sized objects. Overall, the visualization emphasizes the dominance of small instances and underscores the dense, fine-grained nature of aerially captured transmission-line defects.
@@ 4.5 Comparisons With Previous Methods
Table LABEL:tab:model_performance summarizes the overall performance of representative detectors on the CSG-ADCD dataset. The comparison covers lightweight and standard variants of YOLO (v5, v8, 10, 11), Transformer-based models such as RT-DETR and Conditional-DETR, as well as classical baselines including RetinaNet and DETR. Key indicators, including the number of parameters, computational complexity (GFLOPs), Precision, Recall, $\text{mAP}_{50}$, and $\text{mAP}_{50:95}$, are reported to provide a comprehensive view of accuracy–efficiency trade-offs. The best values in each column are highlighted in bold for clarity.
@@
To further highlight performance on small objects and defect-specific categories, Table LABEL:tab:model_performance2 reports additional evaluation metrics. These include Recall and $\text{mAP}_{50}$ for Polluted Insulator (jyzwh) and Damaged Pole (dgss), alongside scale-sensitive measures $\text{AP}_{s}$ and $\text{AP}_{m}$. Such results enable a fine-grained comparison of detection
@@
To further illustrate the trade-off between detection accuracy and computational efficiency, the results reported in Table LABEL:tab:model_performance are visualized in Fig. 9. The plot highlights the relative positions of different detectors in terms of $\text{mAP}_{50}$ versus GFLOPs, enabling an intuitive comparison of accuracy–efficiency balance across lightweight YOLO variants, Transformer-based detectors, and the proposed TinyDef-DETR. Notably, TinyDef-DETR achieves an $\text{mAP}_{50}$ of 0.275 with a computational cost of only 65.3 GFLOPs, striking a favorable balance between accuracy and efficiency compared with both lightweight YOLO models and heavier Transformer-based detectors.
To provide a holistic comparison of different detectors, we visualize their performance across multiple evaluation metrics in the form of a radar chart (Fig. 10). The selected models include YOLOv5m, YOLOv8m, YOLO11m, RT-DETR-R18, RT-DETR-x, and TinyDef-DETR, representing both one-stage and Transformer-based detection paradigms. The evaluation encompasses a comprehensive set of indicators, namely Recall, $\text{mAP}_{50}$, $\text{mAP}_{50:95}$, $\text{AP}_{s}$, $\text{AP}_{m}$, Recall of Polluted Insulator (jyzwh), $\text{mAP}_{50}$ of Polluted Insulator, and $\text{mAP}_{50}$ of Damaged Pole (dgss). For enhanced visual comparability, each metric is normalized to its maximum value.

### v7

@@ Abstract
Transmission lines;Defect detection;UAV imagery;Small Object Detection;
@@ 3.1 Overview
In this study, we employ RT-DETR-r18Zhao et al. (2024) as the baseline detection framework owing to its superior accuracy and robustness in complex scenarios. Compared with traditional DETR models, RT-DETR exhibits significantly faster training convergence and improved deployment efficiency, making it more suitable for real-time and resource-constrained applications such as UAV-based inspection tasks Suvittawat et al. (2025).
RT-DETR incorporates an enhanced encoder–decoder architecture together with refined query interaction mechanisms, while for the backbone network, RT-DETR-r18 adopts a CNN-based design, namely the popular ResNet-18He et al. (2016). From the perspective of real-time performance, employing a CNN architecture for feature extraction further enhances the real-time capability of DETR-like models, thereby improving their practical applicability. As a result, RT-DETR not only accelerates convergence and improves detection precision, but also preserves the global context modeling capability of Transformer-based frameworks(Vaswani et al., 2017).In addition, its lightweight optimization strategies—specifically, the design choice to apply the encoder only to high-level semantic features (S5) with relatively small spatial dimensions—effectively reduce resource consumption. This balances performance and speed by significantly decreasing computational overhead and accelerating inference without compromising accuracy, thereby making RT-DETR particularly suitable for post-flight UAV imagery analysis, where large volumes of data must be processed efficiently under computational and energy constraints(Tian et al., 2025).
@@ 3.4 Cross-Stage Dual-Domain Multi-Scale Attention Module
In transmission line defect detection, most objects are small-scale objects such as cracks, rust spots, or insulator damages. These defects often exhibit weak visual saliency and subtle structural variations, making them difficult to distinguish from complex backgrounds like vegetation, sky, or towers. Conventional convolutional networks are inherently constrained by their limited local receptive fields, and simply enlarging the kernel size to capture global context quickly leads to over-parameterization and substantial computational overhead, while also causing oversmoothing and the loss of fine-grained detailsFinder et al. (2025). Similarly, traditional attention mechanisms that focus solely on either spatial or channel domains fail to provide sufficient enhancement for ambiguous and small-scale objectsZhou et al. (2025). Therefore, a more effective mechanism is required to simultaneously preserve detailed structural information, capture global semantic context, and adaptively emphasize defect-related features.
To improve the effectiveness of transmission lines defect detection, we propose the Cross-Stage Dual-Domain Multi-Scale Attention Module (CSDMAM). Given an input feature map $X\in\mathbb{R}^{B\times C\times H\times W}$, a $1\times 1$ convolution is first applied to adjust the channel dimension. The transformed features are then processed in parallel through multiple specialized branches, and their outputs are aggregated via residual summation before being projected by another $1\times 1$ convolution. This design enables the module to capture diverse feature representations while maintaining computational efficiency. The overall architecture of CSDMAM is illustrated in Figure 5.
Figure 5: The architecture of the proposed Cross-Stage Dual-Domain Multi-Scale Attention Module (CSDMAM).
To effectively capture complementary information at different scales and semantic levels, CSDMAM is composed of four parallel branches: a Large-Receptive-Field Branch, a Local Detail Preservation Branch, a Global Branch based on the Frequency–Spatial–Channel Attention (FSCA) module, and a Residual Aggregation Branch.
@@
To strengthen semantic discrimination and capture long-range contextual information, we design a Global Branch built upon the Frequency–Spatial–Channel Attention (FSCA) module, which jointly exploits frequency, spatial, and channel dependencies. Given the input features $X\in\mathbb{R}^{B\times C\times H\times W}$, a two-dimensional Fourier transform is first applied to obtain their frequency-domain representation:
@@
where $\mathcal{F}$ denotes the discrete Fourier transform (DFT).
Channel-wise weights are then computed via global average pooling followed by a $1\times 1$ convolution, producing $A_{c}\in\mathbb{R}^{B\times C\times 1}$. These weights are used to enhance the frequency representation:
@@
which is subsequently transformed back into the spatial domain using the inverse DFT:
@@
Next, a spatial–channel reweighting map $A_{sc}$ is generated through global pooling and a $1\times 1$ convolution. This map adaptively emphasizes defect-relevant regions and further refines the feature representation, resulting in the final output of the Global Branch:
@@
To effectively integrate the complementary information captured by all branches, we introduce a residual aggregation branch. In this step, the outputs of the large-receptive-field branch, the local branch, and the FSCA branch are combined with the original input through element-wise summation. This residual fusion mechanism not only preserves the original feature information but also enhances gradient flow and stabilizes training. By consolidating global context, fine-grained details, and dual-domain attention responses into a unified representation, the residual aggregation branch ensures comprehensive feature enhancement and precise defect localization. The final aggregated response is formulated as
@@
where $DW$ denotes depth-wise convolution with the specified kernel size.
Finally, to further improve computational efficiency and facilitate gradient propagation, we extend the design into a cross-stage structure. Following the Cross Stage Partial (CSP) strategy, the input is split into two parts: one processed by the multi-branch attention module and the other preserved as an identity mapping. The two parts are then concatenated and fused using a $1\times 1$ convolution to produce the final output:
$Y_{cs}=\text{Conv}\left([\text{DMAM}(X_{\text{branch}}),X_{\text{identity}}]\right).$
By jointly leveraging large receptive fields, local detail preservation, FSCA-based dual-domain enhancement, and cross-stage fusion, CSDMAM achieves a favorable balance between global semantic understanding and precise defect localization. This design significantly improves the detection of small, visually ambiguous defects in UAV inspection imagery.
@@ 3.5 Focaler-Wise-SIoU Loss Function
In transmission line defect detection, most objects are extremely small and exhibit subtle spatial variations, which makes precise bounding box regression essential for reliable detection. Intersection-over-Union (IoU)-based losses are widely used in modern detectors for their geometric interpretability and optimization effectiveness, yet they still suffer from fundamental limitations. The canonical IoU loss, which measures only the overlap ratio between predicted and ground-truth boxes, fails to provide informative gradients when the boxes do not intersect, hindering optimization in challenging cases. Extensions such as GIoU (Rezatofighi et al., 2019), DIoU (Zheng et al., 2020), and CIoU (Zheng et al., 2022) address this issue partially by incorporating geometric constraints (e.g., enclosure area, center distance, aspect ratio), but they remain limited in jointly modeling spatial alignment and shape consistency. More advanced designs, such as SIoU (Gevorgyan, 2022), integrate angular, distance, and shape factors, yet they typically treat all samples equally and fail to account for difficulty imbalance, causing easy samples to dominate the optimization while hard ones—particularly small, low-SNR objects—are underemphasized. Moreover, conventional IoU-based losses lack discriminative power in high-IoU regimes, where sub-pixel localization precision is critical, resulting in slow convergence and degraded performance.
To overcome these challenges, we propose Focaler-Wise-SIoU Loss, a unified regression objective that integrates normalized IoU scaling, comprehensive geometric penalties, and a non-monotonic focal modulation mechanism. The proposed formulation enhances gradient sensitivity in high-IoU regions, improves geometric modeling fidelity, and adaptively reweights samples based on their difficulty, thereby accelerating convergence and significantly improving detection accuracy for small, hard-to-localize objects.
Formally, given a predicted bounding box $B_{p}=(x_{p},y_{p},w_{p},h_{p})$ and a ground-truth box $B_{t}=(x_{t},y_{t},w_{t},h_{t})$, parameterized by their center coordinates and width–height pairs, we first convert them into the corner representation $(x_{1},y_{1},x_{2},y_{2})$. The intersection and union areas are computed to yield the standard IoU metric:
$IoU=\frac{S_{\text{int}}}{S_{\text{union}}}.$ (5)
To improve discriminability in high-IoU regions, we apply a linear normalization known as Focal-IoU scaling, mapping $IoU$ into the interval $[0,1]$ as:
$\hat{IoU}=\mathrm{clip}\left(\frac{|IoU-d|}{u-d},0,1\right),$ (6)
where $d=0$ and $u=0.95$. The base IoU penalty is then defined as:
$L_{IoU}=1-\hat{IoU}$. (7)
On top of this formulation, we integrate the Scylla-IoU (SIoU) penalty, which accounts for angular, distance, and shape discrepancies. The angular component is derived from the displacement vector between box centers $\Delta=(d_{x},d_{y})$:
@@
$\mathrm{Angle}=\sin(2\phi)-2.$ (9)
This term modulates the distance penalty as follows:
$\mathrm{Dist}=2-\exp\left(\mathrm{Angle}\cdot\frac{d_{x}}{w_{\mathrm{box}}}\right)-\exp\left(\mathrm{Angle}\cdot\frac{d_{y}}{h_{\mathrm{box}}}\right),$ (10)
where $(w_{\mathrm{box}},h_{\mathrm{box}})$ denote the width and height of the smallest enclosing box. To further capture shape consistency, we define the shape penalty as:
$\mathrm{Shape}=\left(1-\exp\left(-\frac{|w_{p}-w_{t}|}{\max(w_{p},w_{t})}\right)\right)+\left(1-\exp\left(-\frac{|h_{p}-h_{t}|}{\max(h_{p},h_{t})}\right)\right)^{\theta},$ (11)
where $\theta=4$. The resulting SIoU-based core loss is expressed as:
$L_{SIoU}=L_{IoU}+\frac{\mathrm{Dist}+\mathrm{Shape}}{2}.$ (12)
To adaptively account for sample difficulty, particularly in datasets dominated by small and hard-to-localize objects, we introduce a Wise Modulation mechanism. A relative difficulty factor is defined as:
$\beta=\frac{L_{IoU}}{\mathbb{E}[L_{IoU}]},$ (13)
where $\mathbb{E}[L_{IoU}]$ is an exponential moving average updated with a momentum of $10^{-2}$. Unlike conventional monotonic modulation strategies, which monotonically emphasize harder samples and risk overfitting to outliers, we employ a non-monotonic focal modulation function:
$\gamma(\beta)=\frac{\beta}{\delta\cdot\alpha^{(\beta-\delta)}},$ (14)
where $\delta$ and $\alpha$ are modulation hyperparameters. This formulation selectively amplifies moderately difficult samples while suppressing trivial or excessively noisy cases, contributing to more stable and balanced training dynamics.
Finally, the complete Focaler-Wise-SIoU Loss is defined as:
$L=\left(1-\hat{IoU}+\frac{\mathrm{Dist}+\mathrm{Shape}}{2}\right)\cdot\frac{1-IoU}{\mathbb{E}[1-IoU]}\cdot\left(\frac{1-IoU}{\mathbb{E}[1-IoU]}\right)^{\delta}.$ (15)
This unified formulation preserves the geometric interpretability and simplicity of traditional IoU-based losses while significantly enhancing discriminative capacity in high-IoU regimes, adaptively balancing sample contributions, and stabilizing the optimization process. As a result, the proposed Focaler-Wise-SIoU Loss demonstrates superior robustness and precision in transmission line defect detection, where most objects are small, difficult to localize, and require highly accurate bounding box regression.
@@ 4.1 CSG-ADCD
The CSG-ADCD (China Southern Grid Aerial Defective Component Dataset) is a large-scale, high-resolution UAV-based image dataset specifically developed for the detection of defects in power transmission line components, including insulators, tie wires, and poles. Constructed under authentic inspection scenarios by China Southern Power Grid, the dataset comprises 10,000 aerial images. Each image was manually annotated by a single domain expert to ensure both consistency and high-quality labeling.
In total, the dataset contains 73,448 annotated instances covering nine component states or defect types: Normal Insulator (zcjyz), Polluted Insulator (jyzwh), Damaged Pole (dgss), Missing Tie Wire (zxqs), Loose Tie Wire (zxst), Insulator Flashover (jyzsl), Bird Nest (nw), Broken Insulator (jyzps), and Shattered Post Insulator (zyzsl).
@@
Table 1: Correspondence between labels and category names
No. Label Name
@@
As shown in Fig.6, the vast majority of objects are smaller than 32×32 pixels, placing them within the small-object category. The scatter plot depicts the distribution of annotated instances by pixel width and height after preprocessing, with different categories distinguished by distinct colors in the legend. The strong concentration of points in the lower-left corner highlights the predominance of small and compact objects, consistent with the characteristics of transmission-line inspection imagery. Dashed reference lines at 32×32 and 96×96 pixels mark the boundaries between small, medium, and large regimes, clearly revealing the relative scarcity of medium- and large-sized objects.
Overall, the visualization emphasizes the dominance of small instances and underscores the dense, fine-grained nature of aerially captured transmission-line defects.
@@ 4.5 Comparisons With Previous Methods
Table 3 summarizes the overall performance of representative detectors on the CSG-ADCD dataset. The comparison covers lightweight and standard variants of YOLO (v5, v8, 10, 11), Transformer-based models such as RT-DETR and Conditional-DETR, as well as classical baselines including RetinaNet and DETR. Key indicators, including the number of parameters, computational complexity (GFLOPs), Precision, Recall, $\text{mAP}_{50}$, and $\text{mAP}_{50:95}$, are reported to provide a comprehensive view of accuracy–efficiency trade-offs. The best values in each column are highlighted in bold for clarity.
@@
To further highlight performance on small objects and defect-specific categories, Table 4 reports additional evaluation metrics. These include Recall and $\text{mAP}_{50}$ for Polluted Insulator (jyzwh) and Damaged Pole (dgss), which are the two most frequently occurring defect types in transmission line inspection scenarios and are therefore selected as representative categories for detailed evaluation, alongside scale-sensitive measures $\text{AP}_{s}$ and $\text{AP}_{m}$. Such results enable a fine-grained comparison of detection performance.
@@
A comprehensive illustration of the trade-off between detection accuracy and computational efficiency is presented in Fig. 9, where the results reported in Table 3 are visualized. The plot highlights the relative positions of different detectors in terms of $\text{mAP}{50}$ versus GFLOPs, enabling an intuitive comparison of the accuracy–efficiency balance across lightweight YOLO variants, Transformer-based detectors, and the proposed TinyDef-DETR. Notably, TinyDef-DETR achieves an $\text{mAP}{50}$ of 0.275 with a computational cost of only 65.3 GFLOPs, striking a favorable balance between accuracy and efficiency compared with both lightweight YOLO models and heavier Transformer-based detectors.
Furthermore, Table 4 provides a more detailed view of performance on small objects and defect-specific categories. It reports Recall and $\text{mAP}_{50}$ for Polluted Insulator (jyzwh) and Damaged Pole (dgss) — the two most frequently occurring defect types in transmission line inspection — alongside scale-sensitive metrics $\text{AP}_{s}$ and $\text{AP}_{m}$. These results offer deeper insights into model behavior and enable fine-grained performance comparisons across representative defect categories.
Finally, a holistic comparison of the selected detectors is presented in the form of a radar chart (Fig. 10). The models included — YOLO v5m, YOLO v8m, YOLO 11m, RT-DETR-R18, RT-DETR-x, and TinyDef-DETR — represent both one-stage and Transformer-based detection paradigms. The evaluation encompasses a comprehensive set of indicators, namely Recall, $\text{mAP}{50}$, $\text{mAP}{50:95}$, $\text{AP}s$, $\text{AP}m$, Recall of Polluted Insulator (jyzwh), $\text{mAP}{50}$ of Polluted Insulator, and $\text{mAP}{50}$ of Damaged Pole (dgss). For enhanced visual comparability, each metric is normalized to its maximum value.
@@
4.6 Generalization Experiments for the Proposed Methods
Since small objects dominate our dataset, we further evaluate the proposed method on the VisDrone dataset (Zhu et al., 2022), a widely used UAV-based benchmark for object detection and tracking. VisDrone contains over 26,000 high-resolution images and 263 video sequences captured in diverse urban and suburban scenes, with more than 100,000 annotated objects across 10 common categories. Due to its dense object distribution, large proportion of small targets, and complex real-world conditions, VisDrone provides a challenging platform for evaluating model robustness and generalization.